# DiTCtrl: Exploring Attention Control in Multi-Modal Diffusion Transformer for Tuning-Free Multi-Prompt Longer Video Generation

### v1

@@ 4.2 Quantitative Results
Figure 8: Ablation Component in DiTCtrl. The first and second rows have 98 frames, while the remaining methods generate 105 frames.
@@ 4.3 Ablation Study
4.4 More Applications
Single-prompt Longer Video Generation.
Our method can naturally work on single-prompt longer video generation. As illustrated in Fig. 9, using the prompt “A white SUV drives on a steep dirt road”, our approach successfully generates videos that are more than 12 times longer than the original length, while maintaining consistent motion patterns and environmental coherence.
Video Editing. We show how we use our methods to achieve video editing performance (e.g. “reweight” and “word swap”). The cases are provided in the Appendix. C.
Figure 9: single prompt longer video generation example.
@@ 5 Conclusion
In this paper, we introduce DiTCtrl, a novel, tuning-free method for multi-prompt video generation using the MM-DiT architecture. Our pioneering analysis of MM-DiT’s attention mechanism reveals similarities with the cross/self-attention blocks in UNet-like diffusion models, enabling mask-guided semantic control across prompts. With KV-sharing mechanisms and latent blending strategies, DiTCtrl ensures smooth transitions and consistent object motion between semantic segments, without extra training. We also present MPVBench, the first extensive evaluation framework for multi-prompt video generation, set to advance future research in this field.
@@ Appendix A Implementation Details
sampling num frames 13
@@
Baselines. To demonstrate the effectiveness of our proposed DiTCtrl, we conduct comprehensive qualitative comparisons with state-of-the-art multi-prompt video generation methods [33, 40, 38] and leading commercial solutions Kling [3]. Both FreeNoise and Video-Infinity are built upon the VideoCrafter2 [11] framework. To ensure a fair comparison of base models, we implement FreeNoise [33] as an enhanced baseline by directly incorporating their noise rescheduling strategy into the CogVideoX framework.
@@
$F_{fusion}=F_{fore}*M_{i}+F_{back}*(1-M_{i})$ (9)
Figure 10: Mask-guided KV-sharing details.
Appendix B More Qualitative Results
More results are provided in Fig. 11 and Fig. 12. Our method DiTCtrl can generate multi-prompt videos with good temporal consistency and strong prompt-following capabilities, demonstrating cinematographic-style transitions in depicting the boy’s riding sequence. We also give more qualitative comparisons with state-of-the-art multi-prompt video generation methods [33, 40, 38], our reproduced FreeNoise+DiT, and leading commercial solutions Kling [3]. We show the motion transition case, and background transition case in Fig. 13 and Fig. 14. Our comparative analysis reveals distinct characteristics and limitations of existing approaches. Gen-L-Video [40] suffers from severe temporal jittering and occasional subject disappearance, compromising overall video quality. Video-Infinity [38] and FreeNoise [33] both demonstrate successful scene-level semantic changes but lack physically plausible motion - for instance, in Fig. 13, dark knight appear to be in motion while remaining spatially fixed, which is a limitation inherent to their UNet-based abilities. In contrast, FreeNoise+DiT leverages the DiT architecture’s abilities to achieve more realistic object motion but struggles with semantic transitions, resulting in noticeable discontinuities between segments. Our proposed DiTCtrl method preserves the inherent capabilities of the pre-trained DiT model while addressing these limitations, enabling smooth semantic transitions and maintaining motion coherence throughout the video sequence. More comparison of visualization case and our results are shown in our project page.
Figure 11: More multi-prompt results
Figure 12: More multi-prompt results
Figure 13: Motion and background transition.
Figure 14: Background transition.
Appendix C Applications
@@
C.1 Single-prompt Longer Video Generation
Figure 15: Visualization of single prompt longer video generation.
@@ C.2 Video Editing
Figure 17: Word Swap example of Video Editing.
Appendix D Prompt Generator
We use GPT4 for longer multi-prompt generation, our prompts are shown in Fig. 19. This figure shows the generation process of ”background transition”, and we generate 10 different transition modes (background transition, subject transition, camera transition, style transition, lighting transition, location transition, speed transition, emotion transition, clothing transition, action transition). The detailed prompts will be provided in the the prompts.txt in the attached supplementary.
Appendix E Ablation Study
E.1 Quantitative Results of Components
As shown in Tab. 4, our latent blending strategy (second row) demonstrates superior video consistency compared to isolated clips (first row), as evidenced by higher CSCV scores - our proposed metric for evaluating multi-prompt transition smoothness. Furthermore, our KV-Sharing mechanism further improves the CSCV value, achieving enhanced stability. The mask-guided approach(fourth row) and its unmasked counterpart(third row) report comparable scores, suggesting that the contribution of masking foreground object to overall frame transition smoothness is modest. However, our qualitative analysis in Section E.2 reveals that the mask-guided method yields superior visual results.
Additionally, in our evaluation of motion smoothness, our full method (DiTCtrl) achieves optimal performance. Regarding the Text-Image similarity metric, we observe a slight expected decrease with our approach. This is attributable to our methodology where the latent representation of latter video segments incorporates Keys and Values from preceding segments to maintain consistency. This inherently introduces semantic information from previous segments, marginally reducing the current segment’s alignment with its corresponding text prompt. However, this trade-off is justified as our method achieves stable transitions and effectively conveys both semantic elements, resulting in higher user study scores as shown in Tab. 2.
@@ E.1 Quantitative Results of Components
E.2 Mask-guided Generation Analysis

### v2

@@ 4.2 Qualitative Results
4.3 Quantitative Results
We conduct the automatic evaluation with our MPVBench. From Table 1 one can see that our method achieves the highest CSCV score, demonstrating superior transition handling and overall stability in generation patterns. While FreeNoise ranks second with relatively strong stability, other methods significantly lag behind in this aspect, which is consistent with the t-SNE visualization of CLIP embedding as shown in Fig. 7. In terms of motion smoothness, our approach exhibits superior performance in motion quality and consistency. Regarding Text-Image Similarity metrics, although FreeNoise and Video-Infinity achieve higher scores, this can be attributed to our method’s kv-sharing mechanism, where subsequent video segments inherently learn from preceding semantic content.
As shown in Fig. 6, our design choice allows the road surface to gradually transition to snowy conditions while retaining features from the previous scene. Despite potentially lower text-image alignment scores, it ensures superior semantic continuity in the sequences. In practice, this trade-off doesn’t negatively impact the visual quality in multi-prompt scenarios, as demonstrated by our user study in Table 2.
4.4 Human Evaluation
We invited 28 users to evaluate five models: Gen-L-Video [43], Video-Infinity [41], FreeNoise [36], FreeNoise+DiT and our method. We employ a Forced Ranking Scale, where items are ranked from 1 to 5, with the highest rank receiving a score of 5 and the lowest rank receiving a score of 1. Participants score each method considering overall preference, motion pattern, temporal consistency and text alignment over 16 videos generated by different scenarios. As clearly indicated in Table 2, generated videos from our method significantly outperform other state-of-the-art approaches in all four criteria, demonstrating superior capability in producing videos with natural semantic transitions that better align with human preferences for visual coherence and continuity.
@@ 4.4 Human Evaluation
4.5 More Applications
Single-prompt Longer Video Generation. Our method can naturally work on single-prompt longer video generation. As illustrated in Fig. 8, using the prompt “A white SUV drives on a steep dirt road”, our approach successfully generates videos that are more than 12 times longer than the original length, while maintaining consistent motion patterns and environmental coherence.
Video Editing. We show how we use our methods to achieve video editing performance (“word swap” and “reweight”) in Fig. 9.
  • Word Swap: removing our latent blending strategy of our approach DiTCtrl, we can achieve the video editing performance of Word Swap. Specifically, we just use masked-guided KV-sharing strategy to share keys and values from source prompt $P_{source}$ branch, so that we can synthesize a new video to preserve the original composition while also addressing the content of the new prompt $P_{target}$.
  • Reweight: Similar to prompt-to-prompt [19], through reweighting the specific columns and rows corresponding to specified token (e.g. “pink”) in the MM-DiT’s Text-Video attention and Video-Text attention, we can also achieve the video editing performance of reweight.
Figure 8: Single prompt longer video generation example.
Figure 9: Video editing example.
@@ 4.6 Ablation Study
Figure 10: Visualization of ablation component in DiTCtrl.
@@ 5 Conclusion
In this paper, we introduce DiTCtrl, a novel, tuning-free method for multi-prompt video generation using the MM-DiT architecture. Our pioneering analysis of MM-DiT’s attention mechanism reveals similarities with the cross/self-attention blocks in UNet-like diffusion models, enabling mask-guided semantic control across prompts. With mask-guided kv-sharing mechanism and latent blending strategy, DiTCtrl ensures smooth transitions and consistent object motion between semantic segments, without extra training. We also present MPVBench, a new benchmark with diverse transition types and specialized metrics for assessing multi-prompt transitions.
@@
Acknowledgements. This work is partially supported by the National Natural Science Foundation of China (Grant No. 62306261), and The Shun Hing Institute of Advanced Engineering (SHIAE) Grant (No. 8115074).
@@ Appendix A Implementation Details
video frames 49
latent num frames 13
@@
Baselines. In experiments of our main paper, we comprehensively compare our method with previous state-of-the-art methods, including commercial and open-source techniques. We offer more details of the baselines that we use here:
  • Kling [2]: Kling is leading closed-source commercial solutions developed by Kuaishou Technology. It can generate videos of 6s lengths, but it can only input single-prompt, so we input a single prompt describing long-term temporal changes. We use the Kling1.5 model for our visualization comparison.
  • Gen-L-Video [43]: Gen-L-Video processes long videos as short video clips with temporal overlapping during the denoising process. We use the VideoCrafter2 [11] as the base model.
  • FreeNoise [36]: FreeNoise reschedules the initial noise sequence and conducts temporal attention fusion based on the sliding window for temporal consistency. We use the VideoCrafter2 [11] as the base model.
  • Video-Infinity [41]: Video-Infinity scales up long video generation via distributed inference. We use the VideoCrafter2 [11] as the base model.
  • FreeNoise+DiT: This is an enhanced baseline by directly incorporating FreeNoise’s noise rescheduling strategy into the CogVideoX [48] framework.
For a fair comparison, all baseline methods should be aligned to use the same ratio stride. Since CogVideoX-2B has 13 latent frames, we used overlap frame 6 in our paper which is approximately 1/2 stride of the total frames ($6/13\approx 1/2$). Other baseline methods also use this setting of same stride ratio.
$F_{fusion}=F_{fore}*M_{i}+F_{back}*(1-M_{i})$ (9)
Figure 11: Mask-guided KV-sharing details.
Appendix B Evaluation details
@@
t-SNE Visualization discussion. In the justification for the proposed CSCV metric, which evaluates the transition smoothness, We found that t-SNE visualizations of real videos from existing datasets have similar continuous trajectories due to semantic continuity. Therefore, we just present one representative case, the t-SNE of video embeddings for real videos. The selected real video in Fig.7 of main paper is the classic car video from DAVIS [34]. The car video frames are shown in Fig.12.
We also show more t-SNE visualization of our comparison cases in Fig. 18 and Fig. 19. Even when processing multi-prompt videos, our method generates continuous trajectories that are comparable to those in real videos. This showcases the exceptional transition handling capabilities and overall stability of the videos produced by DiTCtrl.
Figure 12: real video example from DAVIS [34].
Figure 13: t-SNE Visualization of Fig. 18
Figure 14: t-SNE Visualization of Fig. 19
Appendix C More Qualitative Results
More results are provided in Fig. 16 and Fig. 17. Our method DiTCtrl can generate multi-prompt videos with good temporal consistency and strong prompt-following capabilities, demonstrating cinematographic-style transitions in depicting the boy’s riding sequence. We also give more qualitative comparisons with state-of-the-art multi-prompt video generation methods [36, 43, 41], our reproduced FreeNoise+DiT, and leading commercial solutions Kling [2]. We show the motion transition case, and background transition case in Fig. 18 and Fig. 19. Our comparative analysis reveals distinct characteristics and limitations of existing approaches. Gen-L-Video [43] suffers from severe temporal jittering, compromising overall video quality. Video-Infinity [41] and FreeNoise [36] both demonstrate successful scene-level semantic changes but lack physically plausible motion. For instance, in Fig. 18, dark knight appear to be in motion while remaining spatially fixed, which is a limitation inherent to their UNet-based abilities. In contrast, FreeNoise+DiT leverages the DiT architecture’s abilities to achieve more realistic object motion but struggles with semantic transitions, resulting in noticeable discontinuities between segments. Our proposed DiTCtrl method preserves the inherent capabilities of the pre-trained DiT model while addressing these limitations, enabling smooth semantic transitions and maintaining motion coherence throughout the video sequence. More comparison of visualization case and our results are shown in our project page.
Appendix D Applications
@@
D.1 Single-prompt Longer Video Generation
@@ D.2 Video Editing
Appendix E Prompt Generator
In this section, we provide additional information of the prompt generator that is described in our main paper. We use GPT4 for longer multi-prompt generation, our prompts are shown in Fig. 23. This figure shows the generation process of ”background transition”, and we generate 10 different transition modes (background transition, subject transition, camera transition, style transition, lighting transition, location transition, speed transition, emotion transition, clothing transition, action transition).
Appendix F Ablation Study
F.1 Quantitative Results of Components
As shown in Tab. 4, our latent blending strategy (second row) demonstrates superior video consistency compared to isolated clips (first row), as evidenced by higher CSCV scores - our proposed metric for evaluating multi-prompt transition smoothness. Furthermore, our KV-Sharing mechanism further improves the CSCV value, achieving enhanced stability. The mask-guided approach(fourth row) and its unmasked counterpart(third row) report comparable scores, suggesting that the contribution of masking foreground object to overall frame transition smoothness is modest. However, our qualitative analysis in Section F.2 reveals that the mask-guided method yields superior visual results.
Additionally, in our evaluation of motion smoothness, our full method (DiTCtrl) achieves optimal performance. Regarding the Text-Image similarity metric, we observe a slight expected decrease with our approach. This is attributable to our methodology where the latent representation of the latter video segments incorporates keys and values from preceding segments to maintain consistency. This inherently introduces semantic information from previous segments, marginally reducing the current segment’s alignment with its corresponding text prompt. However, this trade-off is justified as our method achieves stable transitions and effectively conveys both semantic elements, resulting in higher user study scores as shown in Tab. 2.
@@ F.1 Quantitative Results of Components
F.2 Mask-guided Generation Analysis
@@
Appendix G Inference Time
We present a comparison of the inference times on a single A100 GPU, with the variation based on the number of prompts (N). For a fair assessment, when 2 prompts are input, each method is tasked to generate approximately 100 frames. When the number of prompts increases to 3, the generation target is set at approximately 150 frames. As depicted in Table 5, our method (without mask) demonstrates competitive efficiency in terms of elapsed time, and also achieves satisfactory video transition effects. When the mask-guided approach is further employed, it yields even more superior visual outcomes. Despite the sixfold increase in runtime, the method remains Pareto optimal.
Gen-L-Video FreeNoise FreeNoise+DiT Video-Infinity Ours(w/o mask) Ours(w/ mask)
N=2 9.1min 6.1min 5.3min 1.2min (2 gpu) 5.3min $\sim$39min
N=3 13.6min 9.2min 10.6min 1.2min (3 gpu) 10.6min $\sim$78min
Table 5: Inference time comparison with the number of prompts N
Figure 16: More multi-prompt results
Figure 17: More multi-prompt results
Figure 18: Motion and background transition.
Figure 19: Background transition.
Figure 20: Visualization of single prompt longer video generation.
Figure 21: Reweighting example of Video Editing.
Figure 22: Word Swap example of Video Editing.